using singularity functions plot the shear and moment diagrams. The magnitude of the moment at point b is 4 kN/m
kN/m В със К 1 2
Answer #1

SKN UKN/m free body Diagram RA= Reaction froce at a Ro= Reaction force at D B C D ykNm im -- Shear force Diagram (KN) . 1.75mTake moment at A=0 Rox4 = 4X6X6 + 4+8x3 Ro= 25kN Equaling vertical forces VRAARD = 8+4X6 RA + 2S = 8+29 RA=7KN shear force EqVD=+74-03°-434-07-8371-370 = -17 KM for point E put >= 6m VE = +756-09-436-07-85:6-37 °+2576-45° Ve = O KN Bending moment Equfor point E put rez Gom ME=726-071 – 426-072+436-9598<6-38 +25T6-4Y ME = 0 KVM